OpenStudy (bamboo888):

Ella had 1/3 left os a wall to paint in her room. She painted 1/4 of what was left to paint as green and 1/2 of it as yellow. What is the area of the region that Ella has NOT YET painted? Length/Width=8ft/12ft A 88ft2 B 16ft2 C 8ft2 D 1ft2 I got C, is this correct?

OpenStudy (bamboo888):

Please help!!!

OpenStudy (bamboo888):

Will medal!

OpenStudy (anonymous):

Total Area = Length x Width = 8 x 12 = 96 ft^2 Area Left = 1/3 of 96 =1/3 x 96 =32 ft^2 Green Area = 1/4 of 32 =1/4 x 32 =8 ft^2 Yellow Area = 1/2 of 32 =1/2 x 32 =16 ft^2 Area not yet painted = Area Left - (Green Area + Yellow Area) = 32 - (8 + 16) = 32 - 24 = 8 ft^2 C So yes, you are correct, well done

OpenStudy (bamboo888):

Thank you!